OpenStudy (anonymous):

please help with the simultaneous equasions y=x^2+15x+43 y=2x+3 just formulae please

OpenStudy (anonymous):

i know that you do 2x+3=x^2+15x+43

OpenStudy (amistre64):

since y=y, y-y=0 y=x^2+15x+43 y=2x+3 x^2+15x+43 = 2x+3 x^2+15x+43-2x-3 = 0

OpenStudy (amistre64):

it looks like if this related to your previous problem, that you added instead of subtracted

OpenStudy (anonymous):

ah yes thank you verry much feel free to ask me any questions

OpenStudy (amistre64):

youre welcome, good luck :)